OpenStudy (anonymous):

sin ^4 x + 2 sin ^2 x cos ^2 x + cos ^4 x

OpenStudy (anonymous):

\(\large sin ^4 x + 2 sin ^2 x cos ^2 x + cos ^4 x=(sin^2x+cos^2x)^2 =(1)^2=1\)

OpenStudy (anonymous):

how come (sin^2 x + cos^2 x) ??

OpenStudy (anonymous):

all you need to do is factor it out... notice the trig expression you gave is of this form: x^2 + 2xy + y^2 and that can be factored to (x + y)^2

OpenStudy (anonymous):

ok